# Security Code Review by Large Language Models

### v2

@@ 4.2. RQ2 Quality Problems in Responses
From the perspective of the four themes, a vast majority of responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below, the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
@@
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to the situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing overly general solutions. An example of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appear in the responses generated by GPT-4. Such information includes the reasons for the absence of security defect in the current code, as well as general suggestions on security measures and review processes. As illustrated in the example below, the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
@@
Compliance: Three types of non-compliance were identified with regard to whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and solutions of detected security defects requested in the prompt. From the example below, we can see that while the potential memory management issues caused by temporary files and directories are described and a solution is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
4.3. RQ3 Factors Influencing LLMs
@@ 4.3.1. Model Stability
To assess the adequacy of our regression model’s fit, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation present in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
4.3.2. Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 6, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (¡.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$¡.0001) and ($\chi^{2}$=18.28, $p$¡.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
@@ 4.3.2. Explanatory Power of Each Factor
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(¿Chi)
TK 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
CMT Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FT Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
ASR -1:0 0.97 2 9.95 0.0069**
@@
AE 0.87 1 4.89 0.0271*
CT 1.01 1 3.06 0.0801
SDT Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
@@
4.3.3. Relationship between Factors and Responses
@@
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 6). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files that contain functional logic, have fewer tokens, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
5. Implications
@@
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (Steenhoek et al., 2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing the accuracy of detection. However, in real-world security reviews, although with the experience and domain knowledge, the reviewer is not guaranteed to consider all potential security defect types in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to effectively focus on the target defect, thereby yielding adverse effects.
@@
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, it appears that annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations there are in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend developers write code comments in a standardized way to clarify the purpose of each comment, thus better support LLMs in code review tasks.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amount of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs to overlook certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and LLama 2. Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
@@ 6. Threats to Validity
Internal Validity: One threat could be data leakage. The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs. Since the training data time of GPT-3.5 and GPT-4 is not clearly disclosed, we cannot avoid this overlap by filtering the data for a specific time period. Moreover, according to the findings by Cao et al. (Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cut-off date. Another threat is the design of prompts. To evoke the best model performance, we design our prompt according to the best practices. We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness. To mitigate this threat, we conducted three repetitive experiments for RQ1. For RQ2, we calculated the average frequency of each problem across these three experiments. In RQ3, we combined the responses from the three experiments to fit the model.
@@
External validity: In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale). In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies. We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
@@
Construct validity: With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure. These models are also not made publically available via API, thereby hindering the reproducibility of this study. Therefore, we explicitly indicate the versions of these two models used in the experiments, and provided all the responses generated by these models (Anonymous, 2024). In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements. Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
7. Conclusions

### v3

@@ 4.2. Quality Problems in Responses (RQ2)
From the perspective of the four themes, most responses contain problems related to Conciseness, accounting for 90.18%. Nearly half of the responses (41.56%) involve Compliance related problems, which indicate the requirements of the task described in the prompt are not met. Moreover, 27.08% of the responses show problems of Understandability due to inappropriate logic, expression, or attribution. Notably, problems related to Correctness, such as deviation from objective facts or common knowledge in the field of security, are encountered less frequently than other themes (14.08%). The typical types of problems within each theme are described in detail below.
Correctness: The responses reveal two types of incorrectness: Incorrect fact (6.25%) and Incorrect concept (10.13%). The responses with Incorrect fact include: incorrect line numbers of the code, function names not matching the source code provided, names not corresponding to the CWE-ID, and the described code content not matching the actual scenario. Incorrect concept indicates a misunderstanding of the concept of a certain CWE-ID or a specific security defect type. As shown in the example below (Example 1), the inaccurate comparison logic described in the response was mistakenly categorized as CWE-682, thus the response is considered to exhibit the Incorrect concept problem.
@@
Understandability: Problems that reduce understandability of responses include Vague statement (21.38%), Incoherent expression (1.19%), Logical contradiction (1.11%), and Groundless speculation (5.26%). Among these, Vague statement is the most prevalent one. It refers to a situation where the description of the detected security defect is not specific enough, failing to pinpoint the defect to a specific location or providing a overly general fix suggestion. An example (Example 2) of Vague statement is shown below, in which GPT-4 did not provide a concrete location or line number for the identified security defect in the response.
@@
Conciseness: In all responses, 11.4% contain Redundant content, 15.12% include information Irrelevant to the task of security defect detection, and 86.38% contain Unnecessary information, all of which are problems related to conciseness of responses. Among them, Unnecessary information, which refers to information that is somewhat related to security code review but does not contribute to detecting security defects, frequently appears in the responses generated by GPT-4. Such information includes the reasons for the absence of security defects in the current code and general suggestions on security measures and review processes. As illustrated in the example below (Example 3), the response indicates that no defect related to CWE-284 exists in the code and further elaborates on the reason. However, this is not the content we required GPT-4 to generate in the prompt, hence is deemed to be Unnecessary information.
@@
Compliance: Three types of non-compliance were identified regarding whether the responses satisfied the requirements in the prompts. Among them, Detection failure is observed in very few responses (0.15%), where GPT-4 explicitly claimed that it could not perform security check in its response or misinterpreted requirements in the prompts for other tasks. Non-compliant output was present in 20.43% of responses, in which GPT-4 fails to include the fixed phrase ‘No security defects are detected in the code’, when no defects are found. Incomplete answer is the most common problem (20.98%) under this theme, referring to cases in which GPT-4’s responses lack the line numbers and fix suggestions of detected security defects requested in the prompt. From the example below (Example 4), we can see that while the potential memory management issues caused by temporary files and directories are described and a fix suggestion is provided, no specific line numbers are given. Hence, it is categorized as an Incomplete answer.
@@
While Kabir et al. (Kabir et al., 2024) also evaluated the quality problems in LLM-generated responses, the application domain of LLMs and the distribution of identified problems in their work differ from ours. Specifically, Kabir et al. requested LLMs to answer programming questions on Stack Overflow, finding that over half (52%) of the responses suffered from correctness issues. In contrast, our study applied LLMs to security code review, revealing a much lower incidence of correctness issues at 14.81%. This indicates that LLMs face different challenges across various software engineering tasks, highlighting the need to optimize and customize LLMs for specific areas of application.
4.3. Factors Influencing LLMs (RQ3)
@@ 4.3.1. Model Stability
To assess the adequacy of the fit of our regression model, we utilized Nagelkerke’s Pseudo $R^{2}$ (Nagelkerke et al., 1991). This yielded a value of 0.158, which is relatively lower than the value of regression analysis in other fields reported in e.g., (Paul et al., 2021; Turzo, 2022). A possible explanation is that although we have considered the randomness of LLM responses during model fitting, the random effects incorporated into our model may not be sufficient to capture all the random variation in the dataset. Nonetheless, through a log-likelihood ratio test using the anova function of the ordinal R package (Christensen and Christensen, 2015), we found that this regression model exhibits significant explanatory power ($LR\,chi2=129.57$, $P_{r}(>chi2)<0.0001$) over a null model. Given that our goal is to construct an inferential regression model rather than a predictive one, this model with a relatively lower $R^{2}$ can still provide valid insights into the relationship between explanatory variables and response variables (Allison, 2014).
@@
Factor (:Reference) Odds Ratio D.F. $\chi^{2}$ Pr(>Chi)
Token 0.92 1 48.07 <.0001absentsuperscript.$0001^{***}< .$
Community Qt:Openstack 1.94 1 19.03 <.0001absentsuperscript.$0001^{***}< .$
FileType Source:Auxiliary 2.82 1 18.28 <.0001absentsuperscript.$0001^{***}< .$
AnnotationisSecurityRelated -1:0 0.97 2 9.95 0.0069**
@@
AuthorExperience 0.87 1 4.89 0.0271*
Commit 1.01 1 3.06 0.0801
SecurityDefectType Buffer Overflow:Race Condition 0.68 10 13.98 0.1742
@@
4.3.2. Explanatory Power of Each Factor
The Wald statistic is employed to evaluate the explanatory power exerted by each factor on the response variable. A larger Wald $\chi^{2}$ with a smaller p-value indicates that the variable has more significant predictive power for the regression model. We calculated the $\chi^{2}$ and p-value of each factor and adopted the commonly used significance level of 0.05 to evaluate their impact. As shown in Table 7, ‘Token’ wields the highest predictive power on the fitted model, with the highest $\chi^{2}$ (48.07) and the lowest p-value (<.0001). Trailing closely behind are ‘Community’ and ‘FileType’ factors, with correspondingly significant contributions ($\chi^{2}$=19.03, $p$<.0001) and ($\chi^{2}$=18.28, $p$<.0001). Adding to this, ‘AnnotationisSecurityRelated’ and ‘AuthorExperience’ displayed less influence but remained noteworthy, with $\chi^{2}$ values of 9.95 and 4.89, respectively. However, the remaining factors, i.e., ‘Commit’, ‘SecurityDefectType’, ‘Complexity’, ‘AnnotationRatio’, and ‘AnnotationHasCode’ are deemed not to largely influence the performance of LLMs in detecting security defects in code reviews, as their p-values fall below the 0.05 threshold. Notably, ‘SecurityDefectType’ has a relatively high $\chi^{2}$ value (13.98), which could be attributed to the fact that this factor is a nominal variable with multiple categories of security defects, thus automatically allocated a high number of degrees of freedom. As a result, even if this factor is not significant, its relative contribution ($\chi^{2}$) remains elevated.
4.3.3. Relationships between Factors and Responses
@@
The odds ratio (OR) is commonly used to measure the relationships between explanatory and response variables (Paul et al., 2021; Thongtanunam et al., 2017; Setia, 2016). In this study, the OR represents the odds of LLM generating a better response in security code review versus not generating a better response. For categorical explanatory variables, the OR compares each category to a reference category. For interval variables, the OR reflects the effect of a one-unit increase in the explanatory variable. We utilized the summary function of the rms package (Harrell Jr et al., 2017) to calculate the OR for each factor (see Table 7). The OR of ‘FileType’ exceeds 1, being 2.82, while the OR of ‘Token’ and ‘AuthorExperience’ are both less than 1, at 0.92 and 0.87, respectively. This suggests that the LLM is more likely to detect security defects in code files that have fewer tokens, contain functional logic, and were written by developers with less involvement in the project. Furthermore, the OR of ‘Community’, which is greater than 1, indicates that compared to OpenStack, the LLM performs better in security code reviews from Qt. In addition, for ‘AnnotationisSecurityRelated’, when annotations in code files contain descriptions related to the target security defect (labelled as 1), the OR is 2.69, indicating that with related security annotations, the LLM is more likely to identify the target security defect. Conversely, when annotations contain descriptions related to non-target security defects (labelled as -1), the OR is 0.97. This suggests that the LLM may be disrupted by annotations and less likely to detect the target security defect.
5. Discussion
A multi-layered review strategy is suggested to apply LLMs to security code review. As the results of RQ1 indicate, the existing general-purpose LLMs still fall significantly short of reaching the effectiveness of manual security code review. However, LLMs can serve as auxiliary tools to assist reviewers. The findings of RQ3 suggest that LLMs are more adept at identifying relatively simple security defects in shorter, logic-centric code written by developers with less project experience. We recommend a multi-layered review strategy to fully leverage the capability of LLMs - reviewers can first use LLMs to conduct an initial automatic review of those short and simple commits for security defects, then review both the submitted code and the LLM’s detection results to improve efficiency. This would allow reviewers to concentrate on an in-depth review of those complex commits, including lengthy code written by more experienced developers, thereby reducing the risk of missing sophisticated security defects. Additionally, where LLMs currently show limitations, these complex commits should be more heavily incorporated into datasets used for fine-tuning LLMs to improve their effectiveness in security code review.
@@ 5. Discussion
Providing a task scope is crucial for enhancing the performance of LLMs. Our findings indicate that incorporating a CWE list into prompts significantly improves the performance of various LLMs in security defect detection, which aligns with the results of Steenhoek et al. (Steenhoek et al., 2024). Providing a CWE list can effectively narrow the problem domain, guiding LLMs to focus on key areas, thereby increasing detection accuracy. However, in real-world security reviews, although with the experience and domain knowledge, The reviewers lack prior knowledge of all potential types of security defects that may be present in the code. Therefore, constructing a comprehensive range of CWEs covering commonly encountered security defect types is necessary. It should also be noted that an overly exhaustive scope can diminish the accuracy of LLM responses. In Bakhshandeh et al.’s work (Bakhshandeh et al., 2023), providing a list of all 75 CWEs in their prompts resulted in a decrease in the accuracy of security defect detection and location by the LLMs. This drop in accuracy may be attributed to the overly lengthy and detailed list of CWEs provided, which interferes with the LLM’s capability to focus effectively on the target defect, yielding adverse effects.
@@
Using LLMs in security code review requires developers to standardize their code commenting practices. According to RQ3, annotations containing information related to non-target security defects, which were often already resolved by developers, negatively impact the LLM’s capability in security code review. This suggests that LLMs could mistakenly interpret these resolved security defects as existing defects. Moreover, although the p-value of ‘AnnotationRatio’ (0.21) is more than 0.05, we still consider its OR value (0.56) to be somewhat informative. The more annotations in the code, the more challenging it is for LLMs to detect security defects, reflecting the disruptive effect and poor quality of annotations written by developers. Hence, we recommend that developers write code comments in a standardized way to clarify the purpose of each comment, thus better-supporting LLMs in code review tasks.
@@
With the development of LLMs, while expanding its context window, it is also necessary to ensure that LLMs can accurately capture details from long inputs. The emergence of LLMs that support large amounts of tokens has alleviated the challenge of inputting code context information in applying LLMs to code-related tasks. However, as evidenced by the results of RQ3, when dealing with code that contains more tokens, it is more difficult for LLMs to identify target security defects. This is because many security defects are introduced by small code snippets, and longer inputs may distract LLMs from overlooking certain details, thereby failing to detect security defects introduced by small code snippets. Additionally, according to RQ1, despite GPT-4 Turbo has a context of 128K tokens, its performance in detecting security defects is noticeably inferior to that of smaller-context LLMs like GPT-4, GPT-3.5, and Llama 2. Therefore, how to ensure the accuracy of security defect detection when expanding the LLMs’ context window is a promising direction.
A key challenge in applying LLMs to security code review is ensuring comprehensive detection of security defects while maintaining consistent results. The results from RQ1.4 indicate that LLMs with stronger capabilities to detect security defects tend to exhibit poorer consistency in their responses. This may be because the outputs generated by LLMs with higher randomness are more diverse, making them more likely to include content related to target security defects compared to those generated by LLMs with lower randomness. Both detection capability and the stability of detection results are crucial for the practical application of LLMs in security code review. However, current studies related to LLMs in vulnerability detection have largely overlooked the importance of the stability of results. Therefore, for researchers of LLMs, exploring approaches to balance detection capabilities with the randomness inherent in LLMs represents a valuable direction for future improvements.
@@ 6. Threats to Validity
Internal Validity. One threat could be data leakage. The dataset used in our study covers two large-scale open-source projects, which may overlap with the data used for training LLMs. Since the training data cutoff dates for GPT-3.5 and GPT-4 are not disclosed and the dataset we used (Yu et al., 2023) predates the release of some of the LLMs we selected, we cannot avoid the overlap between the dataset we used and the training data of LLMs by filtering the data for a specific period of time. Moreover, according to the findings by Cao et al. (Cao et al., 2024), contaminated data does not always affect the results, and the model sometimes even performs better on data after the models’ cutoff date. Another threat is the design of prompts. We design our prompt according to best practices to evoke the best model performance. We selected words that are as unbiased, not suggestive, and unambiguous as possible and iteratively revised the prompts to avoid individual words from impacting the responses of LLMs during our preliminary exploration. In addition, we only included code from the single vulnerable file rather than the whole patchset in the prompt, which may affect the performance of LLMs due to the lack of contextual information. To specifically analyze its impact, we designed prompt $\textbf{P}_{cot}$, in which we instructed LLMs to generate code context according to the commit message before detecting security defects. Lastly, LLMs may generate varied responses for the same prompts due to their inherent randomness. To mitigate this threat, we conducted three repetitive experiments for RQ1. For RQ2, we calculated the average frequency of each problem across these three experiments. In RQ3, we combined the responses from the three experiments to fit the model.
@@
Conclusion Validity. In LLM selection, we tried to encompass a diverse range of popular LLMs (i.e., open-source and closed-source, large-scale and small-scale). In dataset collection, we used a dataset of code reviews collected from real-world open-source projects that are widely used in code review-related studies. We acknowledge that the dataset may not fully represent industrial projects, and projects from proprietary software development are needed to increase the generalizability of the results.
@@
External Validity. With the active development of LLMs, both GPT-3.5 and GPT-4 adopted in our work are continually updated without public disclosure. These models are also not made publically available via API, thereby hindering the reproducibility of this study. Therefore, we explicitly indicate the versions of these two models used in the experiments and provide all the responses generated by these models (Yu et al., 2024). In addition, the data labeling and extraction processes in this study were manually conducted, which may introduce subjective judgements. Thus, we adopted an open coding and constant comparative approach, conducting pilot experiments before all formal manual tasks to ensure a consensus among multiple authors.
7. Conclusions and Future Work